OpenStudy (anonymous):

#1: If you want to make 0.5 L of a 0.01 M solution of bromine (Br2) in water, how much bromine would you need? A. 0.8 grams B. 0.8 moles C. 1.6 grams D. 1.6 moles ***my answer: C. 1.6 grams is that right? @aaronq :)

OpenStudy (anonymous):

#2: Which of the following statements is the correct definition of molarity? A. number of moles of solvent/liter solution B. number of grams of solvent/liter solution C. number of moles of solute/liter solution D. number of grams of solute/liter solution ***my answer: C is that right? :/ not 100% sure...

OpenStudy (aaronq):

1 is wrong :(

OpenStudy (aaronq):

2 is right

OpenStudy (anonymous):

darn :( idk #1 then :( would it be D. 1.6 moles? not grams? and yay for #2 tho! :D

OpenStudy (aaronq):

\[M =\frac{ \frac{ 1.6g }{ 160g/mol } }{ 0.5L }\] = 0.02 M

OpenStudy (anonymous):

#3: You add 0.5 moles of NaCl to water to make a 250 mL solution. What is the molarity of this solution? A. 1.0 M B. 2.0 M C. 2.5 M D. 4.0 M ***my answer: B. 2.0 M is that right?

OpenStudy (aaronq):

1.6/0.5 = 0.32 M so not that either

OpenStudy (aaronq):

3 is good

OpenStudy (anonymous):

okay, so i think it's grams tho right? so would that make it A. 0.8 grams?

OpenStudy (aaronq):

yes, it's 0.8 g buuuut you should be able to work that out

OpenStudy (anonymous):

#4: You have a 2.0 M solution of LiF. You had added 4.00 moles of LiF to a solvent to make this solution. What is the volume of this solution? A. 2 L B. 3 L C. 6 L D. 8 L ***my answer: A. 2L is that right? :)

OpenStudy (anonymous):

yeah haha i know! lol so idk how to do that then :( can you please show me?

OpenStudy (aaronq):

just use the molarity formula and try different values knowing the volume you're using

OpenStudy (aaronq):

4 is good btw

OpenStudy (anonymous):

#5: You have a solution made of 300 grams of ethanol (C2H5OH) and 500 grams of water. What is the mole fraction of ethanol in this solution? A. 0.19 M B. 0.81 M C. 0.19 D. 0.81 ***my answer: C. 0.19 is that right? :)

OpenStudy (anonymous):

ohh okay i see :) and yay for 4!! :) what about 5? :/

OpenStudy (aaronq):

5 is good (Y)!

OpenStudy (anonymous):

omg! yay! =) thank you!!! :)

OpenStudy (aaronq):

no problem !